OpenStudy (anonymous):

Help please... I'm terrible at math can someone tell me the value of X and how to find it? This is the problem 5^(x-1)=2^x

OpenStudy (anonymous):

\[5^{x-1}=2^x\] take log of both side \[\log 5^{x-1}=\log 2^x\] pull out the exponent \[{x-1}\log 5=xlog 2\]

OpenStudy (anonymous):

xlog 5 -log5=xlog2 xlog 5-xlog2=log5

OpenStudy (anonymous):

x(log 5 - log 2)=log 5

OpenStudy (anonymous):

Thank you so much... but... what is X?

OpenStudy (anonymous):

divide both side by (log 5- log 2) then you would have x by itself

OpenStudy (anonymous):

Oh wow! Thanks! I totally love you!

OpenStudy (anonymous):

then you may award me a medal or fan me

OpenStudy (anonymous):

How do I do that? This is literally my first time...

OpenStudy (shadowfiend):

Click good answer on the right of their answer :) You can hover over their avatar to become a fan!

OpenStudy (anonymous):

thanks to both of you!